PREDICTION OF FUNDAMENTAL VIBRATIONAL FREQUENCIES AND INFRARED INTENSITIES: A BENCHMARK STUDY

\underline{JUANA VÁZQUEZ}, MICHAEL E. HARDING, JOHN F. STANTON, Institute for Theoretical Chemistry, Department of Chemistry and Biochemistry, University of Texas at Austin, Austin, TX 78712.

In this work we investigate the performance of second-order vibrational perturbation theory (VPT2) using force fields computed at the fc-CCSD(T) level in conjunction with different double-, triple-, and quadruple-ζ basis sets^{a} for the prediction of fundamental vibrational frequencies and infrared intensities. A benchmark study comprising more than thirty small and medium sized molecules illustrates the accuracy and limitations of the presented scheme.

aAtomic natural orbital (ANOY, Y=0,1,2) [J. Almlöf and P. R. Taylor, J. Chem. Phys. 86, 4070 (1987)] and correlation-consistent (cc-pVXZ, X=D,T,Q) [T. H. Dunning, Jr., J. Chem. Phys. 90, 1007 (1989)] basis sets.